# Zak-OTFS for Spread Spectrum Communication

### v1

@@ Abstract
An attractive feature of spread spectrum technologies such as code division multiple access (CDMA) is that it is harder to intercept or jam signals, and this feature was lost when orthogonal frequency domain modulation prevailed over CDMA in wireless standards. Legacy spread carrier waveforms are not matched to delay and Doppler shifts characteristic of 6G wireless environments, and this makes equalization very challenging. Zak-OTFS (orthogonal time frequency space) modulation is a communication framework that parameterizes the wireless channel in the delay-Doppler (DD) domain, where the parameters map directly to physical attributes of the scatterers that comprise the scattering environment. As a consequence, the channel can be efficiently acquired and equalized. The Zak-OTFS carrier is a pulse in the DD domain, and the Zak transform converts it to a pulse train modulated by a tone (pulsone) in the time domain. The pulsone waveform is localized rather than spread, and it suffers from high peak-to-average power ratio (PAPR).
We describe how to transform Zak-OTFS into a spread spectrum communication system, where the spread carrier waveforms have low PAPR and are matched to the delay and Doppler characteristics of the wireless channel. This transformation is realized by a unitary transform that is a generalization of the discrete affine Fourier transform. The transform maps a pulsone to a time domain waveform which yields a constant amplitude zero auto-correlation (CAZAC) sequence after sampling. The family of CAZAC sequences includes the Zadoff-Chu sequences incorporated in LTE and 5G-NR standards. We describe the end-to-end time-domain transceiver signal processing, comprising channel estimation and data demodulation, for the proposed system. We quantify system performance through bit error rate (BER) simulations using a six-path Vehicular-A channel model, showing that the proposed system achieves similar uncoded BER as pulsone-based Zak-OTFS, where the PAPR of each spread carrier waveform is only $3.58$ dB.
@@
6G, Zak-OTFS, Peak-to-Average Power Ratio, Zadoff-Chu Sequences
@@ I Introduction
Zak-OTFS (orthogonal time frequency space modulation) [1, 2] is an attractive framework for communication, radar sensing, and integrated sensing and communications that processes signals in the delay-Doppler (DD) domain. The Zak-OTFS carrier waveform is a pulse in the DD domain, formally a quasi-periodic localized function with specific periods along delay and Doppler. When the channel delay spread is less than the delay period, and the channel Doppler spread is less than the Doppler period, the Zak-OTFS Input/Output (I/O) relation changes slowly and predictably, and it is not subject to fading [1, 2]. The response to a single Zak-OTFS carrier provides an image of the the scattering environment, and can be used to predict the I/O response to all other carriers. This property makes the Zak-OTFS carrier an ideal radar waveform that can enable integrated sensing and communication in 6G wireless [2].
The Zak-OTFS carrier is realized in the time domain as a pulse train modulated by a tone (pulsone), which suffers from high peak-to-average power ratio (PAPR) [3]. It is possible to reduce PAPR significantly by applying a unitary spreading filter to pulsones in the DD domain [3]. We apply a unitary transform to pulsones in the time domain with the property that the resulting spread carrier waveforms, when sampled, give constant amplitude zero auto-correlation (CAZAC) sequences [4, 5, 6]. Since the transform is unitary, it enables spread carrier communication at the Nyquist rate in the presence of mobility and delay spread. The family of CAZAC sequences includes Zadoff-Chu (ZC) sequences that are widely used in LTE [7] and 5G-NR [8] wireless systems as reference signals [9], uplink random access preambles [10, 11, 12] and downlink synchronization signals [13]. The larger CAZAC family shares the auto-correlation and low PAPR properties [6, 5, 14] that make ZC sequences attractive.
Our contribution in this paper is as follows. We introduce a discrete generalization of the fractional Fourier transform that is unitary so that correlation properties of pulsones transfer to spread carriers without change. We then describe the end-to-end time-domain transceiver signal processing, comprising channel estimation and data demodulation, for the proposed system. We quantify system performance through bit error rate (BER) simulations using a six-path Vehicular-A channel model [15], showing that the proposed system achieves similar uncoded BER as pulsone-based Zak-OTFS, where the PAPR of each spread carrier waveform is only $3.58$ dB.
@@ II Preliminaries: Zak-OTFS via Pulsones
Consider an $M\times N$ delay-Doppler (DD) grid with $M$ delay bins (indexed by $k$) and $N$ Doppler bins (indexed by $l$), corresponding to width $\tau_{p}$ along delay and width $\nu_{p}$ along Doppler with $\tau_{p}\nu_{p}=1$. The basis element corresponding to the $(k_{0},l_{0})$th DD bin, where $0\leq k_{0}\leq(M-1),~{}0\leq l_{0}\leq(N-1)$, is $\delta[k-k_{0}]\delta[l-~{}l_{0}]$. The corresponding discrete time realization via the inverse discrete Zak transform is [16]:
@@ III Spread Carrier Waveforms in Zak-OTFS
The GDAFT in Definition 1 maps the discrete time point pulsone given by (1) localized in the discrete DD domain at $(k_{0},l_{0})$ to the CAZAC sequence:
@@
IV Temporal Spread Spectrum Processing
@@ IV-B Time Domain System Model
We describe a general system model applicable to both pulsone-based and spread carrier-based Zak-OTFS. After mounting the information symbols on either basis, the time domain signal is transmitted through an effective channel which encompasses the effects of the linear time-varying physical channel and transmit & receive pulse shaping [2, Eq. (7)]. The received time domain signal is given by:
@@ V Numerical Results
We simulate the system model in (6) for the pulsone and spread carrier systems with parameters $M=17,N=19$, Doppler period $\nu_{p}=30$ KHz. The effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ in (4) corresponds to sampling the continuous channel [1, 2, 3]:
@@
where $BW=M\nu_{p},T=N\tau_{p}$. The physical channel in (V), $\mathbf{h}_{\mathrm{phy}}(\tau,\nu)=\sum_{i=1}^{P}h_{i}\delta(\tau-\tau_{i})\delta(\nu-\nu_{i})$, corresponds to a 3GPP-compliant $P=6$ path Vehicular-A (Veh-A) channel model [15] with significant mobility and delay spread, whose power-delay profile is shown in Table I. The Doppler of each path is simulated as $\nu_{i}=\nu_{\max}\cos(\theta_{i})$, with $\theta_{i}$ uniformly distributed in $[-\pi,\pi)$ and $\nu_{\max}=815$ Hz denoting the maximum channel Doppler spread. Note that the delay and Doppler of each path in the model is fractional in nature. For the spread carrier basis, we choose the GDAFT parameters $A=3,B=5,C=7$ such that the crystallization condition in Lemma 2 holds. The transmitted data symbols $\mathbf{x}_{d}$ in (6) are chosen from $4$-QAM (quadrature amplitude modulation).
TABLE I: Power-delay profile of Veh-A channel model
@@
(a) Pulsone basis element.
(b) Spread carrier basis element.
Figure 1: Time-domain waveforms of pulsone and spread carrier basis elements corresponding to $(k_{0},l_{0})=((M+1)/2,(N+1)/2)$. The former has a PAPR of $15.1$ dB, whereas the latter has a PAPR of $3.58$ dB. Note that only the real part of a portion of the entire waveform is shown.
@@ V-A Peak-to-Average Power Ratio (PAPR) Comparison
Fig. 1 plots the time-domain waveforms of the pulsone and spread carrier basis elements with a sinc transmit pulse shaping filter. Due to the constant amplitude property of CAZACs, the latter has $11.5$ dB lower PAPR than the former.
@@ V-B Channel Estimation Performance
Fig. 2 plots the normalized mean squared error (NMSE) in channel estimation via the cross-ambiguity expression in (8) for pulsone- and spread carrier-based Zak-OTFS. We consider two cases - (i) equal signal-to-noise ratio (SNR) in the pilot and data stages in (6), and (ii) with $40$ dB pilot SNR. The NMSE of the former reduces with increasing pilot SNR and approaches the NMSE of the latter at $40$ dB SNR. In both cases, the NMSEs match for pulsones and spread carriers, i.e., the systems are equivalent. Note that channel estimates only need to be accurate enough (not perfect) for data detection.
@@ V-C Data Detection Performance
Fig. 3 plots the bit error rate (BER) curves for pulsone- and spread carrier-based Zak-OTFS for both cases described in Section V-B as well as perfect channel knowledge. The BER with $40$ dB pilot SNR is equivalent to that with perfect channel knowledge. The BER curve for equal pilot and data SNR approaches the other curves with increasing SNR. In all cases, the BERs match for pulsones and spread carriers.
@@ VI Conclusion
In this paper, we proposed a spread spectrum communication system using Zak-OTFS by transforming high-PAPR pulsones to low-PAPR CAZAC sequences via a unitary transformation. We demonstrated the equivalence of both systems via bit error rate simulations with a six-path Vehicular-A channel model. Future work will further explore the application of the proposed spread spectrum Zak-OTFS in secure communication, radar sensing, and integrated sensing & communication.

### v3

# Zak-OTFS with Spread Carrier Waveforms

@@ Abstract
Zak-OTFS (orthogonal time frequency space) modulation is a communication framework that parameterizes the wireless channel in the delay-Doppler (DD) domain, where the parameters map directly to physical attributes of the scatterers that comprise the scattering environment. As a consequence, the channel can be efficiently acquired and equalized. The Zak-OTFS carrier is a pulse in the DD domain, and the Zak transform converts it to a pulse train modulated by a tone (pulsone) in the time domain. The pulsone waveform is localized rather than spread, and it suffers from high peak-to-average power ratio (PAPR). We describe how to transform the orthonormal basis of Zak-OTFS pulsones into an orthonormal basis of spread carrier waveforms with low PAPR (only $6.58$ dB) that support communication in the presence of mobility and delay spread. This transformation is realized by a unitary transform based on the discrete affine Fourier transform. Unlike other spread modulations that achieve low PAPR by spreading information across a wider bandwidth (thus reducing the spectral efficiency), the proposed spread carrier-based Zak-OTFS achieves full spectral efficiency like pulsone-based Zak-OTFS, with $5.6$ dB lower PAPR per basis element. We demonstrate uncoded bit error rate (BER) similar to pulsone-based Zak-OTFS, and improved BER performance over competing methods based on OFDM and OTFS in high mobility & delay spread environments.
@@
6G, Zak-OTFS, Peak-to-Average Power Ratio, Spread Carrier Communication
@@ I Introduction
Zak-OTFS (orthogonal time frequency space modulation) [1, 2] is an attractive framework for communication, radar sensing, and integrated sensing and communications that processes signals in the delay-Doppler (DD) domain. The Zak-OTFS carrier waveform is a pulse in the DD domain, formally a quasi-periodic localized function with specific periods along delay and Doppler. When the channel delay spread is less than the delay period, and the channel Doppler spread is less than the Doppler period, the Zak-OTFS Input/Output (I/O) relation changes slowly and predictably, and it is not subject to fading [1, 2]. The response to a single Zak-OTFS carrier provides an image of the the scattering environment, and can be used to predict the I/O response to all other carriers. The image of the scattering environment can support sensing in systems such as passive radars that integrate sensing and communications [2].
TABLE I: Comparison with prior work. $\alpha$ denotes the ratio of total vs used subcarriers in DFT-spread-OFDM ($\alpha>1$).
Method Spectral Eff. PAPR Predictable?
Spread Zak-OTFS (Ours) 1 6.6 dB
Pulsone Zak-OTFS [1, 2] 1 12.2 dB
DFT-spread-OFDM [3] $\nicefrac{{1}}{{\alpha}}$ 6–8 dB $\times$
OTFS [4, 5, 6] 1 10–12 dB $\times$
OFDM 1 10–12 dB $\times$
The Zak-OTFS carrier is realized in the time domain as a pulse train modulated by a tone (pulsone), which suffers from high peak-to-average power ratio (PAPR) [7]. It is possible to reduce PAPR significantly by applying a unitary spreading filter to pulsones in the DD domain [7]. We apply a unitary transform to pulsones in the time domain to generate spread carrier waveforms with constant amplitude zero auto-correlation (CAZAC [^1] Examples of CAZAC sequences are Zadoff-Chu (ZC) sequences [8] that are used for synchronization and random access in LTE and 5G-NR.[9, 10] properties. The proposed unitary transform is based on a generalization of the discrete affine Fourier transform [11].
Since the transform is unitary, the proposed method achieves full spectral efficiency ($BWT$ information symbols in $BWT$ channel uses) like pulsone-based Zak-OTFS, with $5.6$ dB lower PAPR per basis element. In contrast, existing spread modulations like DFT-spread-OFDM [3] achieve low PAPR by spreading information across a wider bandwidth; thus have spectral efficiency inversely proportional to the spreading factor. Table I places our contributions in the context of prior work. In Section V, through bit error rate (BER) simulations using a six-path Vehicular-A channel model, we demonstrate uncoded $4$-QAM222quadrature amplitude modulation BER similar to pulsone-based Zak-OTFS, and improved BER performance over competing methods based on OFDM, DFT-spread-OFDM [3] and OTFS [4, 5, 6] in high mobility & delay spread environments.
@@ II Preliminaries: Zak-OTFS via Pulsones
Consider an $M\times N$ delay-Doppler (DD) grid [^3] OFDM equivalence: $M$ is number of subcarriers, $N$ is number of symbols, $\nu_{p}$ is subcarrier spacing. There is no cyclic prefix in Zak-OTFS. with $M$ delay bins (indexed by $k$) and $N$ Doppler bins (indexed by $l$), corresponding to width $\tau_{p}$ along delay and width $\nu_{p}$ along Doppler with $\tau_{p}\nu_{p}=1$. The basis element corresponding to the $(k_{0},l_{0})$th DD bin, where $0\leq k_{0}\leq(M-1),~{}0\leq l_{0}\leq(N-1)$, is $\delta[k-k_{0}]\delta[l-~{}l_{0}]$. The corresponding discrete time realization [^4] Superscript $(\cdot)^{(\mathrm{p})}$ denotes pulsone. via the inverse discrete Zak transform is [12]:
@@ III Spread Carrier Waveforms in Zak-OTFS
Proof:
We may represent the GDAFT by an ${MN\times MN}$ matrix $\mathbf{U}$, with the $(n,m)$th element of $\mathbf{U}$ given by $\mathbf{U}[n,m]=\frac{1}{\sqrt{MN}}e^{\frac{j2\pi}{MN}(An^{2}+Bnm+Cm^{2})}$. Therefore, $\mathbf{U}^{\mathsf{H}}\mathbf{U}$ is given by:
$(\mathbf{U}^{\mathsf{H}}\mathbf{U})[m_{1},m_{2}]$ $=\!\frac{1}{MN}\!\sum_{n=0}^{MN-1}\!e^{\frac{j2\pi}{MN}(Bn+C(m_{2}+m_{1}))(m_{2}-m_{1})}$
$=\delta[m_{1}=m_{2}],$
where the final expression follows from Identity 1 and the fact that $0\leq m_{1},m_{2}\leq(MN-1)$. Hence, $\mathbf{U}^{\mathsf{H}}\mathbf{U}$ is the $MN\times MN$ identity matrix, i.e., the GDAFT is a unitary transform. ∎
@@
The GDAFT in Definition 1 maps the discrete time point pulsone given by (1) localized in the discrete DD domain at $(k_{0},l_{0})$ to a spread carrier CAZAC sequence [^5] Superscript $(\cdot)^{(\mathrm{c})}$ denotes spread carrier.:
@@
Proof:
From Lemma 1, we know that the GDAFT is a unitary transform. Therefore, the GDAFT maps orthonormal bases to orthonormal bases and preserves inner products. Since pulsones form an orthonormal basis [1, 2], the GDAFT maps the pulsone basis to another orthonormal basis formed by CAZACs of the form given in Theorem 1. ∎
IV Time Domain Spread Carrier Processing
@@ IV-B Time Domain System Model
We describe a general system model applicable to both pulsone-based and spread carrier-based Zak-OTFS. After mounting the information symbols on either basis, the time domain signal is transmitted through an effective channel [^6] The effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $BW$ and in Doppler with time $T$. which encompasses the effects of the linear time-varying physical channel and transmit & receive pulse shaping [2, Eq. (7)]. The received time domain signal is given by:
@@ IV-C Channel Estimation (Pilot Stage)
In practice, the GDAFT parameters $A,B,C$ must be chosen such that the crystallization condition in Lemma 2 holds. We illustrate the choice of parameters $A,B,C$ via an example.
[…]
In practice, grid search is performed to determine $A,B,C$. Note that prior knowledge of channel support $\mathcal{S}$ does not need to be exact, and estimates of $k_{\min},k_{\max},l_{\min},l_{\max}$, e.g., from the maximum expected delay to the cell edge and maximum speed of objects, can be used to approximate $\mathcal{S}$.
(a) $A=3,B=5,C=7$.
(b) $A=2,B=5,C=7$.
Figure 1: Example illustrating choice of GDAFT parameters.
@@ V Numerical Results
We simulate the system model in (6) for both pulsone-based and spread carrier-based Zak-OTFS. The effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ in (4) is generated by sampling the continuous effective channel as [1, 2, 7]:
@@
where $BW=M\nu_{p},T=N\tau_{p}$. The physical channel in (V), $\mathbf{h}_{\mathrm{phy}}(\tau,\nu)=\sum_{i=1}^{P}h_{i}\delta(\tau-\tau_{i})\delta(\nu-\nu_{i})$, corresponds to a 3GPP-compliant $P=6$ path Vehicular-A (Veh-A) channel model [16] with significant mobility and delay spread, whose power-delay profile is shown in Table II. The Doppler of each path is simulated as $\nu_{i}=\nu_{\max}\cos(\theta_{i})$, with $\theta_{i}$ uniformly distributed in $[-\pi,\pi)$ and $\nu_{\max}=815$ Hz denoting the maximum channel Doppler spread [^9] Our channel model considers fractional delay and Doppler shifts, which is representative of real propagation environments. The path delays in Table II are non-integer multiples of the delay resolution $\nicefrac{{1}}{{BW}}$. The Doppler shifts $\nu_{i}=\nu_{\max}\cos(\theta_{i})$ are also non-integer multiples of the Doppler resolution $\nicefrac{{1}}{{T}}$ since $\cos(\theta_{i})$ is continuous valued.. We simulate both narrowband and wideband system regimes with parameters: $M=17,N=19$, $BW=0.51$ MHz (narrowband) and $M=83,N=13$, $BW=19.9$ MHz (wideband). For the spread carrier basis, we choose the GDAFT parameters $A=3,B=5,C=7$ such that the crystallization condition in Lemma 2 holds. The transmitted data symbols $\mathbf{x}_{d}$ in (6) are chosen from an uncoded $4$-QAM constellation.
TABLE II: Power-delay profile of Veh-A channel model
@@
Figure 2: PAPR of pulsone- and spread carrier-based Zak-OTFS.
@@ V-A Peak-to-Average Power Ratio (PAPR) Comparison
Fig. 2 plots the complementary cumulative distribution function (CCDF) of the PAPR for pulsone- and spread carrier-based Zak-OTFS for system parameters: $M=17,N=19$, $BW=0.51$ MHz and an oversampling factor of $4$. Due to the CAZAC property of the spread carrier basis (see Theorem 1), each spread carrier basis element has $5.6$ dB lower PAPR than each pulsone basis element. When modulated by information symbols, the spread carrier data signal has $7.83$ dB PAPR compared to $7.95$ dB PAPR for the pulsone data signal.
@@ V-B Channel Estimation Performance
Fig. 3 plots the normalized mean squared error (NMSE) in channel estimation via (8) for pulsone- and spread carrier-based Zak-OTFS for system parameters: $M=17,N=19$, $BW=0.51$ MHz. We consider two scenarios: (i) equal signal-to-noise ratio (SNR) in the pilot and data stages [^10] The transmit power remains same for pilot and data frames in this case. in (6), and (ii) with $20$ dB pilot SNR. The NMSE of the former reduces with increasing pilot SNR and approaches the NMSE of the latter at $20$ dB SNR. In both scenarios, the NMSEs match for pulsones and spread carriers, i.e., the systems are equivalent. We note that for data detection, we only require channel estimates to be accurate enough (but not perfect).
@@ V-C Data Detection Performance
Fig. 4 plots the bit error rate (BER) curves for pulsone- and spread carrier-based Zak-OTFS for system parameters: $M=17,N=19$, $BW=0.51$ MHz and three scenarios: (i) perfect channel knowledge, (ii) equal pilot and data SNR, and (iii) $20$ dB pilot SNR. The BER with $20$ dB pilot SNR diverges from the perfect channel knowledge curve at higher SNRs, matching the equal pilot and data SNR curve at $20$ dB SNR and saturating beyond $20$ dB SNR. In all three scenarios, the BER curves match for pulsones and spread carriers, thus demonstrating the equivalence of pulsone-based and the proposed spread carrier-based Zak-OTFS systems.
V-D Comparison with Existing Approaches
Fig. 5 plots the BER curves for system parameters: $M=17,N=19$, $BW=0.51$ MHz assuming perfect channel knowledge across five systems: pulsone-based Zak-OTFS, spread carrier-based Zak-OTFS, OTFS, OFDM and DFT-spread-OFDM. To ensure frame consistency between the latter three and Zak-OTFS, we choose the subcarrier spacing as $\nu_{p}$, number of subcarriers as $M$, number of symbols per frame as $N$. The cyclic prefix length is $\lceil BW\tau_{\max}\rceil\nicefrac{{1}}{{BW}}$ for maximum delay spread $\tau_{\max}$ of the channel in Table II. Since DFT-spread-OFDM only encodes information on a fraction of the total number of available subcarriers, say $L$ out of the total $M$, for fair comparison we ensure that information is modulated on only $LN$ (randomly chosen) symbols out of the total $MN$ symbols per frame in Zak-OTFS, OTFS and OFDM. Fig. 5 corresponds to the choice $L=9$, for a spectral efficiency of $\nicefrac{{L}}{{M}}=\nicefrac{{9}}{{17}}$. The BER curves demonstrate the value of delay-Doppler signaling with non-fading modulations like Zak-OTFS in doubly-selective channels [1, 2]. The OFDM-based time-frequency modulations have poor BER due to channel fading. OTFS yields significantly better performance than both OFDM systems, and the performance improves further at $15$ dB SNR with Zak-OTFS due to the greater predictability of the Zak-OTFS I/O relation. Moreover, the BER performance of pulsone- and spread carrier-based Zak-OTFS is the same.
Figure 5: Bit error rate curves comparing the performance of the proposed spread Zak-OTFS scheme with existing approaches.
V-E Narrowband vs Wideband Performance
Fig. 6 plots the BER curves for pulsone- and spread carrier-based Zak-OTFS for both narrowband and wideband system parameters assuming perfect channel knowledge. The wideband system parameters are chosen to approximate 5G-NR numerology: $N=13$ (5G-NR has $12-14$ symbols per slot), $\nu_{p}=240$ kHz (5G-NR subcarrier carrier spacing) and $M=83$ for $BW=19.9$ MHz. The BER performance remains similar for narrowband and wideband systems; hence, our proposed method can be utilized in both regimes.
Figure 6: Bit error rate curves showing similar data detection performance for narrowband and wideband systems.
@@ VI Conclusion
In this paper, we proposed a spread carrier implementation of Zak-OTFS by transforming high-PAPR pulsones to low-PAPR spread carrier waveforms via a generalization of the discrete affine Fourier transform. We showed that the proposed system achieves full spectral efficiency like Zak-OTFS with $5.6$ dB lower PAPR per basis element, and outperforms competing methods based on OFDM and OTFS. Future work will explore applications of the proposed method to security, radar sensing, and integrated sensing & communication.